# Towards efficient quantum algorithms for diffusion probability models

### v1

@@
Towards efficient quantum algorithms for diffusion probability models
Yunfei Wang Equal contribution Joint Center for Quantum Information and Computer Science, NIST/University of Maryland, College Park, MD 20742, USA Maryland Center for Fundamental Physics, Department of Physics, University of Maryland, College Park, MD 20742, USA    Ruoxi Jiang Equal contribution Department of Computer Science, University of Chicago Chicago, IL 60637    Yingda Fan Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Xiaowei Jia Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Jens Eisert jense@zedat.fu-berlin.de Dahlem Center for Complex Quantum Systems, Freie Universität Berlin, 14195 Berlin, Germany    Junyu Liu junyuliu@pitt.edu Department of Computer Science, The University of Pittsburgh, Pittsburgh, PA 15260, USA    Jin-Peng Liu liujinpeng@tsinghua.edu.cn Yau Mathematical Sciences Center and Department of Mathematics, Tsinghua University, Beijing 100084, China Yanqi Lake Beijing Institute of Mathematical Sciences and Applications, Beijing 100407, China
A diffusion probabilistic model (DPM) is a generative model renowned for its ability to produce high-quality outputs in tasks such as image and audio generation. However, training DPMs on large, high-dimensional datasets—such as high-resolution images or audio—incurs significant computational, energy, and hardware costs. In this work, we introduce efficient quantum algorithms for implementing DPMs through various quantum ODE solvers. These algorithms highlight the potential of quantum Carleman linearization for diverse mathematical structures, leveraging state-of-the-art quantum linear system solvers (QLSS) or linear combination of Hamiltonian simulations (LCHS). Specifically, we focus on two approaches: DPM-solver-$k$ which employs exact $k$-th order derivatives to compute a polynomial approximation of $\epsilon_{\theta}(x_{\lambda},\lambda)$; and UniPC which uses finite difference of $\epsilon_{\theta}(x_{\lambda},\lambda)$ at different points $(x_{s_{m}},\lambda_{s_{m}})$ to approximate higher-order derivatives. As such, this work represents one of the most direct and pragmatic applications of quantum algorithms to large-scale machine learning models, presumably talking substantial steps towards demonstrating the practical utility of quantum computing.
@@ I Introduction
Diffusion models, in the context of machine learning and, more specifically, deep learning, represent a class of generative models that have garnered significant attention for their novel approach to modeling complex data distributions [1, 2, 3, 3, 4]. Unlike traditional models of machine learning, which often rely on explicit likelihood estimation or adversarial training, diffusion models generate data through a gradual, iterative process inspired by the physical concept of diffusion. They work by initially introducing noise a data points and then learning how to reverse this process, effectively denoising the data step-by-step. This framework allows them to generate high-quality samples, often rivaling or surpassing the performance of other generative models, such as generative adversarial networks (GANs) [5] and variational autoencoders (VAEs) [6]. The appeal of diffusion models lies in their stability during training and their ability to model intricate, high-dimensional data distributions without requiring complex adversarial setups or delicate balancing of training dynamics. Diffusion probabilistic models (DPM), as we call them, are families of generative model renowned for their ability to produce high-quality out- puts in complex tasks such as image and audio generation.
Turning to quantum approaches, quantum computing is widely regarded as one of the most promising alternatives to von Neumann architectures in the post-Moore era [7]. By leveraging unique quantum features such as superposition and entanglement, quantum computing is anticipated to surpass classical counterparts in specific tasks. These include factoring [8], database search [9], simulating complex quantum systems [10], sampling [11], and solving problems in linear algebra [12, 13]. Although we are currently still in what can be called the noisy intermediate-scale quantum (NISQ) era, the concept of a megaquop machine offers a vision of a future where quantum computing achieves remarkable reliability and scalability. Such a machine, with an error rate per logical gate on the order of $10^{-6}$, would surpass the limitations of classical, NISQ, or analog quantum devices. It would have the capacity to execute quantum circuits involving around $100$ logical qubits and a circuit depth of approximately $10,000$—capabilities far beyond the reach of current technology.
Figure 1: Sketch of classical (1.) diffusion probability models (DPM) and the quantum algorithms proposed here for quantum analogs (2.-4.). We pursue two approaches: This is on the one hand Carlemann linearization for DPM solvers (2.), on the other Carleman linearization for UniPC (3.). For both approaches, quantum ODE solvers are instrumental in the step before measurement. QLSS stands for quantum linear system solvers, LCHS for linear combination of Hamiltonian simulation.
@@ II Results
In this section, we lay out our specific results, both concerning rigorous statements that heuristic numerical statements that are meant to provide further evidence for the functioning of the approach. Substantial more detail on all steps will be presented in the supplemental material. We will now explain what steps we take towards establishing a fresh direction for demonstrating the utility of quantum computers in machine learning tasks. Specifically, we present novel and efficient quantum algorithms for implementing diffusion probabilistic models (DPMs) using a variety of ordinary differential equation (ODE) solvers. These algorithms leverage recent advancements in quantum ODE solvers and highlight the potential of quantum Carleman linearization for addressing a broad range of mathematical structures and challenges. Our exploration focuses on two distinct approaches.
@@
These algorithms underscore the potential of quantum Carleman linearization in addressing diverse mathematical structures. Moreover, as we advance toward the era of megaquop machines, this work stands out as one of the most direct and practical applications of quantum algorithms to large-scale machine learning models. It marks, so we hope, a significant milestone in showcasing the tangible utility of quantum computing in real-world applications.
@@ II.1.1 Carleman linearization for DPM-solvers
In this expression $\lambda\coloneqq\log(\alpha_{t}/\sigma_{t})$ is what is called the log-SNR variable. For a given a set of time steps $\{t_{i}\}_{i=0}^{M}$ with $t_{0}=T$ and $t_{M}=0$, with $h_{i}=\lambda_{t_{i}}-\lambda_{t_{i-1}}$, one makes use of the $(k-1)$-st-order Taylor approximation in order to develop what is called the DPM-solver-$k$ in Ref. [23]. This actually takes the form
@@
    $\begin{split}\delta\hat{Y}&=\hat{Y}(t_{i})-\hat{Y}(t_{i-1})=A\hat{Y}(t_{i-1})+b\leavevmode\nobreak\ ,\\ &\text{ for }\quad\hat{y}_{1}(t_{M})=\hat{y}_{1}(0)=x_{\mathrm{in}}(0)\leavevmode\nobreak\ ,\end{split}$ (4)
@@ II.1.2 Carleman linearization for UniPC
    $\begin{split}\delta Y(t_{i})=\leavevmode\nobreak\ &A^{(0)}Y^{(0)}(t_{i})+b^{(0)}\\ &+A^{(1)}Y^{(1)}(t_{i})+b^{(1)}\\ &+\ldots+A^{(p-1)}Y^{(p-1)}(t_{i})+b^{(p-1)}\leavevmode\nobreak\ ,\end{split}$ (7)
@@
    $\begin{split}\delta Z(t_{i})=BZ(t_{i})&+A^{(0)_{c}}Y^{(0)}(t_{i})+b^{(0)_{c}}\\ &+A^{(1)_{c}}Y^{(1)}(t_{i})+b^{(1)_{c}}\\ &+\ldots+A^{(p)_{c}}Y^{(p)}(t_{i})+b^{(p)_{c}}\leavevmode\nobreak\ ,\end{split}$ (8)
@@ II.2 End-to-end quantum ODE solver
$\bm{\mathsf{M}^{c}}\bm{\mathsf{Z}}=\bm{\mathsf{b}^{c}}\leavevmode\nobreak\ ,$ (10)
@@
Besides the QLSS approach, the linear combination of Hamiltonian simulation (LCHS) strategy [33, 34] provides an alternative method for solving the linear ODEs in Eqs. (9) and (10) by representing the solution to the homogeneous equation as a linear combination of Hamiltonian simulations. When a source term (e.g., $\bm{\mathsf{b}^{c}}$ in Eq. (10) and the right-hand side of Eq. (9)) is present, Duhamel’s principle can be applied [33]. LCHS achieves gate efficiency by combining linear combinations of unitaries (LCU) [35] with existing Hamiltonian simulation techniques. The method involves truncating the infinite integral to a finite interval $[-K,K]$ and discretizing it using the $M+1$ grid points. Compared to the standard QLSS approach, LCHS directly implements the time evolution operator, significantly reducing the number of state preparation oracles required.
@@ II.3 Theorems
$\mathcal{O}(J\kappa\cdot\mathrm{polylog}(N/\epsilon))\leavevmode\nobreak\ ,$ (11)
@@
$\mathcal{O}(pJ\kappa\cdot\mathrm{polylog}(N/\epsilon))\leavevmode\nobreak\ ,$ (12)
@@ II.4 Numerical analysis and experiments
Above, we present present performance guarantees of the quantum algorithms introduced. It is a core contribution of this work, however, to also make the point that the approach taken is reasonably applicable to real-world data. This is a valuable contribution, as the proven separations for machine learning problems are for highly structured and artificial data [17, 18, 19]. In this section, we perform extensive numerical experiments to show that for real-world datasets, the conditions of the above theorems are commonly satisfied for natural data and that the processes considered are sufficiently diffusive in a precise sense.
@@
where $a_{i}(t)\coloneqq{\lambda_{i}(t)}/{\max_{i,t}|\lambda_{i}(t)|}$. By construction, $P(t)$ grows unbounded when negative eigenvalues dominate and decays towards zero when positive eigenvalues prevail. As shown in Fig. 2 2., across different dimensionalities of the state variables, the inference process follows a consistent pattern: $P(t)$ steadily decreases as $t$ decreases, implying the emergence of large eigenvalues along the generative process, reinforcing the existence of dissipative modes, as it is required in the assumptions of the theorems.
@@ III Discussion
At a broader level, we aim for our work to help address a key bottleneck in exploring the potential of quantum computing in machine learning. While quantum computers have demonstrated advantages for certain highly structured problems, there is a prevailing view that they may offer limited benefits for tasks in machine learning, which typically involve less structured data and more noise [21]. This work, suggests, however, that there may be significant potential for quantum computers in performing inference tasks or generative tasks of this nature. Considering that inference costs might be more than training costs in many machine learning applications, our proposed approach could pave the way for a practical example of quantum advantage in machine learning industry, while also providing a pathway to making quantum computers more useful for solving industrially relevant problems.
@@ Acknowledgements
J. E. has been supported by the BMBF (Hybrid++, DAQC, QSolid), the QuantERA (HQCC), the BMWK (EniQmA), the DFG (CRC 183), the Quantum Flagship (PasQuans2, Millenion), the Munich Quantum Valley, Berlin Quantum, and the European Research Council (DebuQC). J.-P. L. acknowledges support from Tsinghua University and Beijing Institute of Mathematical Sciences and Applications.

### v2

# Towards efficient quantum algorithms for diffusion probabilistic models

@@
Abstract
A diffusion probabilistic model (DPM) is a generative model renowned for its ability to produce high-quality outputs in tasks such as image and audio generation. However, training DPMs on large, high-dimensional datasets such as high-resolution images or audio incurs significant computational, energy, and hardware costs. In this work, we introduce efficient quantum algorithms for implementing DPMs through various quantum ODE solvers. These algorithms highlight the potential of quantum Carleman linearization for diverse mathematical structures, leveraging state-of-the-art quantum linear system solvers (QLSS) or linear combination of Hamiltonian simulations (LCHS). Specifically, we focus on two approaches: DPM-solver-$k$ which employs exact $k$-th order derivatives to compute a polynomial approximation of $\epsilon_{\theta}(x_{\lambda},\lambda)$; and UniPC which uses finite difference of $\epsilon_{\theta}(x_{\lambda},\lambda)$ at different points $(x_{s_{m}},\lambda_{s_{m}})$ to approximate higher-order derivatives. As such, this work represents one of the most direct and pragmatic applications of quantum algorithms to large-scale machine learning models, presumably taking substantial steps towards demonstrating the practical utility of quantum computing.
@@ I Introduction
Diffusion models, in the context of machine learning and, more specifically, deep learning, represent a class of generative models that have garnered significant attention for their novel approach to modeling complex data distributions [1, 2, 3, 3, 4]. Unlike traditional models of machine learning, which often rely on explicit likelihood estimation or adversarial training, diffusion models generate data through a gradual, iterative process inspired by the physical concept of diffusion. They work by initially introducing noise at data points and then learning how to reverse this process, effectively denoising the data step-by-step. This framework allows them to generate high-quality samples, often rivaling or surpassing the performance of other generative models, such as generative adversarial networks (GANs) [5] and variational autoencoders (VAEs) [6]. The appeal of diffusion models lies in their stability during training and their ability to model intricate, high-dimensional data distributions without requiring complex adversarial setups or delicate balancing of training dynamics. Diffusion probabilistic models (DPM), as we call them, are families of generative models renowned for their ability to produce high-quality out- puts in complex tasks such as image and audio generation.
Turning to quantum approaches, quantum computing is widely regarded as one of the most promising alternatives to von Neumann architectures in the post-Moore era [7]. By leveraging unique quantum features such as superposition and entanglement, quantum computing is anticipated to surpass classical counterparts in specific tasks. These include factoring [8], database search [9], simulating complex quantum systems [10], sampling [11], and solving problems in linear algebra [12, 13]. Although we are currently still in what can be called the noisy intermediate-scale quantum (NISQ) era, the concept of a megaquop machine [14] offers a vision of a future of fault-tolerant application-scale quantum (FASQ) computing [15] in which quantum computing achieves remarkable reliability and scalability. Such a megaquop machine, with an error rate per logical gate on the order of $10^{-6}$, would surpass the limitations of classical, NISQ, or analog quantum devices. It would have the capacity to execute quantum circuits involving around $100$ logical qubits and a circuit depth of approximately 10,00010,000—capabilities far beyond the reach of current technology.
Figure 1: Sketch of classical (1.) diffusion probability models (DPM) and the quantum algorithms proposed here for quantum analogs (2.-4.). We pursue two approaches: This is on the one hand Carleman linearization for DPM solvers (2.), on the other Carleman linearization for UniPC (3.). For both approaches, quantum ODE solvers are instrumental in the step before measurement. QLSS stands for quantum linear system solvers, LCHS for linear combination of Hamiltonian simulation.
@@ II Results
In this section, we lay out our specific results, both concerning rigorous statements that heuristic numerical statements that are meant to provide further evidence for the functioning of the approach. Substantially more detail on all steps will be presented in the supplemental material. We will now explain what steps we take towards establishing a fresh direction for demonstrating the utility of quantum computers in machine learning tasks. Specifically, we present novel and efficient quantum algorithms for implementing diffusion probabilistic models (DPMs) using a variety of ordinary differential equation (ODE) solvers. These algorithms leverage recent advancements in quantum ODE solvers and highlight the potential of quantum Carleman linearization for addressing a broad range of mathematical structures and challenges. Our exploration focuses on two distinct approaches.
@@
These algorithms underscore the potential of quantum Carleman linearization in addressing diverse mathematical structures. Moreover, as we advance toward the era of the era of FASQ [15], this work stands out as one of the most direct and practical applications of quantum algorithms to large-scale machine learning models. It marks, so we hope, a significant milestone in showcasing the tangible utility of quantum computing in real-world applications.
@@ II.1.1 Carleman linearization for DPM-solvers
In this expression $\lambda\coloneqq\log(\alpha_{t}/\sigma_{t})$ is what is called the log-SNR variable. For a given a set of time steps $\{t_{i}\}_{i=0}^{M}$ with $t_{0}=T$ and $t_{M}=0$, with $h_{i}=\lambda_{t_{i}}-\lambda_{t_{i-1}}$, one makes use of the (k1)(k-1)-st-order Taylor approximation in order to develop what is called the DPM-solver-$k$ in Ref. [25]. This actually takes the form
@@
    $\begin{split}\delta\hat{Y}&=\hat{Y}(t_{i})-\hat{Y}(t_{i-1})=A\hat{Y}(t_{i-1})+b\penalty 10000\ ,\\ &\text{ for }\quad\hat{y}_{1}(t_{M})=\hat{y}_{1}(0)=x_{\mathrm{in}}(0)\penalty 10000\ ,\end{split}$ (4)
@@ II.1.2 Carleman linearization for UniPC
    $\begin{split}\delta Y(t_{i})=\penalty 10000\ &A^{(0)}Y^{(0)}(t_{i})+b^{(0)}\\ &+A^{(1)}Y^{(1)}(t_{i})+b^{(1)}\\ &+\ldots+A^{(p-1)}Y^{(p-1)}(t_{i})+b^{(p-1)}\penalty 10000\ ,\end{split}$ (7)
@@
    $\begin{split}\delta Z(t_{i})=BZ(t_{i})&+A^{(0)_{c}}Y^{(0)}(t_{i})+b^{(0)_{c}}\\ &+A^{(1)_{c}}Y^{(1)}(t_{i})+b^{(1)_{c}}\\ &+\ldots+A^{(p)_{c}}Y^{(p)}(t_{i})+b^{(p)_{c}}\penalty 10000\ ,\end{split}$ (8)
@@
A practical limitation of the quantum Carleman linearization lies in the truncation of the polynomial hierarchy, which introduces a systematic truncation error that is difficult to control analytically. The contribution of the neglected higher-order terms generally depends on the nonlinearity strength and on how strongly probability amplitudes populate higher moments during evolution [29]. In complex diffusion models, this error may grow over time or under strong coupling, and we do not yet have a rigorous bound for it. We emphasize this as an open technical limitation rather than a hidden assumption.
That said, for many physically relevant regimes—particularly when $N$ is large and the diffusion dynamics remains near a stable manifold—the effect of higher-order monomials is expected to be suppressed. Empirically, for smooth drifts and moderate noise levels, low-order truncations reproduce accurate dynamics because higher-order correlations decay rapidly or remain small. Thus, although the truncation error is not strictly controllable, it can remain practically negligible within the accuracy relevant for the applications studied here. A full characterization of this error scaling and its dependence on truncation order $N$ and diffusion strength, and truncation order is left for future work.
@@ II.2 End-to-end quantum ODE solver
$\bm{\mathsf{M}^{c}}\bm{\mathsf{Z}}=\bm{\mathsf{b}^{c}}\penalty 10000\ ,$ (10)
@@
Besides the QLSS approach, the linear combination of Hamiltonian simulation (LCHS) strategy [36, 37] provides an alternative method for solving the linear ODEs in Eqs. (9) and (10) by representing the solution to the homogeneous equation as a linear combination of Hamiltonian simulations. When a source term (e.g., $\bm{\mathsf{b}^{c}}$ in Eq. (10) and the right-hand side of Eq. (9)) is present, Duhamel’s principle can be applied [36]. LCHS achieves gate efficiency by combining linear combinations of unitaries (LCU) [38] with existing Hamiltonian simulation techniques. The method involves truncating the infinite integral to a finite interval [K,K][-K,K] and discretizing it using the $M+1$ grid points. Compared to the standard QLSS approach, LCHS directly implements the time evolution operator, significantly reducing the number of state preparation oracles required.
@@ II.3 Theorems
$\mathcal{O}(J\kappa\cdot\mathrm{polylog}(N/\epsilon))\penalty 10000\ ,$ (11)
@@
$\mathcal{O}(pJ\kappa\cdot\mathrm{polylog}(N/\epsilon))\penalty 10000\ ,$ (12)
@@ II.4 Numerical analysis and experiments
Above, we present the performance guarantees of the quantum algorithms introduced. It is a core contribution of this work, however, to also make the point that the approach taken is reasonably applicable to real-world data. This is a valuable contribution, as the proven separations for machine learning problems are for highly structured and artificial data [19, 20, 21]. In this section, we perform extensive numerical experiments to show that for real-world datasets, the conditions of the above theorems are commonly satisfied for natural data and that the processes considered are sufficiently diffusive in a precise sense.
@@
where $a_{i}(t)\coloneqq{\lambda_{i}(t)}/{\max_{i,t}|\lambda_{i}(t)|}$. By construction, $t\mapsto P(t)$ grows unboundedly when negative eigenvalues dominate and decays towards zero when positive eigenvalues prevail. As shown in Fig. 2 2., across different dimensionalities of the state variables, the inference process follows a consistent pattern: $P(t)$ steadily decreases as $t$ decreases, implying the emergence of large eigenvalues along the generative process, reinforcing the existence of dissipative modes, as it is required in the assumptions of the theorems.
@@ III Discussion
At a broader level, we aim for our work to help address a key bottleneck in exploring the potential of quantum computing in machine learning. While quantum computers have demonstrated advantages for certain highly structured problems, there is a prevailing view that they may offer limited benefits for tasks in machine learning, which typically involve less structured data and more noise [23]. This work, suggests, however, that there may be significant potential for quantum computers in performing inference tasks or generative tasks of this nature. What is more, the present ansatz suggests that the quantum advantage is not only present for highly engineered instances, but in fact robust for an entire class of meaningful instances [15]. Considering that inference costs might be more than training costs in many machine learning applications, our proposed approach could pave the way for a practical example of quantum advantage in machine learning industry, while also providing a pathway to making quantum computers more useful for solving industrially relevant problems.
@@ Acknowledgements
J. E. has been supported by the BMFTR (Hybrid++, DAQC, QSolid), the QuantERA (HQCC), the BMWK (EniQmA), the DFG (CRC 183), the Quantum Flagship (PasQuans2, Millenion), the Munich Quantum Valley, Berlin Quantum, and the European Research Council (DebuQC). J.-P. L. acknowledges support from Tsinghua University and Beijing Institute of Mathematical Sciences and Applications.